# On second order conditions for singular optimal control of port-Hamiltonian systems

### v2

@@ 2.2 Optimality conditions for nonlinear ordinary pH systems with control bounds
with real numbers $\underline{u}_{i}\leq\overline{u}_{i},$ for each $i=1,\dots,m$. For the problem with control bounds in the port-Hamiltonian optimal control problem (2.1), the Goh conditions of Theorem 2.1 specialize as described next.
@@
for some pair $i,j=1,\ldots,m,$ it must necessary hold that
$p^{T}[g_{i},g_{j}]=0\quad\text{ and }\quad\ell_{j}^{\prime}\,g_{i}-\ell_{i}^{\prime}\,g_{j}=0.$
In view of (2.4)-(2.5), the results in Theorem 2.2 with control constraints take the following form.
@@
Under the settings of Theorem 2.3, one gets that
$p^{T}[f_{0},[f_{0},f_{i}]]+\sum_{j\in S}u_{j}^{*}p^{T}[f_{j},[f_{0},f_{i}]]$
$+\sum_{j\in\underline{B}}\underline{u}_{j}p^{T}[f_{j},[f_{0},f_{i}]]+\sum_{j\in\overline{B}}\overline{u}_{j}p^{T}[f_{j},[f_{0},f_{i}]]=0.$
holds on singular arcs, where the expression of $w_{ij}\coloneqq p^{T}[f_{0},[f_{0},f_{i}]]$ is given by (2.5) above, $i\in S$ and $S$, $\underline{B}$ and $\overline{B}$ denote the set of singular, lower-bang and upper-bang component of $u^{*}$ In other words, Theorem A.4 in Appendix A holds.
@@
See Theorem A.4 in Appendix A. ∎
Example 2.5.
[…]
the Lie brackets $[f_{i},f_{j}]$ are given by
$[f_{i},f_{j}]=\begin{bmatrix}0\\ 0\\ B_{j}(q)^{\top}\frac{\partial^{2}\mathcal{E}}{\partial p^{2}}B_{i}(q)-B_{i}(q)^{\top}\frac{\partial^{2}\mathcal{E}}{\partial p^{2}}B_{j}(q)\end{bmatrix}.$
These vanish identically for any sufficiently regular $\mathcal{E}$. In the current framework, one has $\mathcal{E}{\partial p^{2}}=M^{-1},$ for $M$ being symmetric positive definite, so Goh conditions (1.10) hold trivially. One gets the following expression for the second-order Lie brackets
$[f_{j},[f_{0},f_{i}]]=\begin{bmatrix}B_{i}^{\prime}M^{-1}B_{j}+B_{j}^{\prime}M^{-1}B_{i}\\ 0\end{bmatrix}\\ +\begin{bmatrix}\frac{\partial D}{\partial p}(M^{-1}B_{i})B_{j}+\frac{\partial^{2}D}{\partial p^{2}}(M^{-1}p,B_{i})B_{j}\\ 0\end{bmatrix},$
@@ 2.3 Optimality conditions for nonlinear pH descriptor systems
In this section we extend the results from the previous two subsections to a special class of pH descriptor systems. For general descriptor systems optimality condition have been derived in [20], see [22] for a detailed exposition. But again these do not exploit the particular structure.
It has been shown in [7] that by an appropriate regularization procedure, using transformations, derivatives, and feedback designs, a general descriptor system can always be transformed to a so-called index-one or strangeness-free system, see [22].
Then for a system of the form (1.1)–(1.4), after appropriate coordinate changes, we may assume that the system has the structure
@@
$\qquad\left(\begin{bmatrix}J_{11}&J_{12}\\ -J_{12}^{T}&J_{22}\end{bmatrix}(x)-\begin{bmatrix}R_{11}&R_{12}\\ R_{12}^{T}&R_{22}\end{bmatrix}(x)\right)\begin{bmatrix}Q_{11}&Q_{12}\\ Q_{21}&Q_{22}\end{bmatrix}\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix}$
@@
Observe that the condition $Q^{T}E=E^{T}Q\geq 0$ implies that $Q_{12}=Q_{21}^{T}=0$ and $Q_{11}\geq 0$. In the following, to keep the presentation simple, we also assume that $G_{2}(x)=0$. If this is not the case then the following reduction procedure will lead to a system with a feed-through term for which an extra transformation is necessary, see [29]. The assumption that the system is of index at most one means that from the algebraic equation
$0=$ $[(J_{22}(x_{1},x_{2})-R_{22}(x_{1},x_{2}))Q_{21}$
$+(-J_{12}^{T}(x_{1},x_{2})-R_{12}^{T}(x_{1},x_{2}))Q_{11}x_{1}+[(J_{22}(x_{1},x_{2})-R_{22}(x_{1},x_{2}))Q_{22}x_{2},$
the variable $x_{2}$ can be expressed via the Implicit Function Theorem as a function of $x_{1}$ and inserted in the coefficients. Due to the structure of $G(x)$, $x_{2}$ contributes to the cost functional only indirectly as argument of $G$. Thus, we can first solve the optimal control problem for $x_{1}$ and obtain $x_{2}$ in a post-processing step. This means that the dynamics of the system is described by the ordinary pH system
@@
3 Optimality conditions for linear pH descriptor systems
@@ 3.1 Optimality conditions for linear pH systems without control constraints
  • $J\in\mathbb{R}^{n,n}$ is skew-symmetric,
  • $R\in\mathbb{R}^{n,n}$ is symmetric positive semidefinite,
  • $Q\in\mathbb{R}^{n,n}$ is symmetric positive definite,
  • $G=\begin{bmatrix}g_{1}&\ldots&g_{m}\end{bmatrix}\in\mathbb{R}^{n,m}$ has full rank $m\leq n,$
@@
Note that more general linear port-Hamiltonian formulations are possible, we follow here the formulation considered in [10, 11, 39].
@@
  • $Y\in\mathbb{R}^{n,n}$ is symmetric and positive semidefinite,
  • $N=\begin{bmatrix}n_{1}&\ldots&n_{m}\end{bmatrix}\in\mathbb{R}^{m,m}$,
@@
$[f_{i},f_{j}]=\begin{bmatrix}0\\ \vdots\\ 0\\ n_{j}^{T}G^{T}Qg_{i}-n_{i}^{T}G^{T}Qg_{j}\end{bmatrix}\quad i,j=1,\dots,m,$
@@
$\begin{bmatrix}p\\ 1\end{bmatrix}^{T}[f_{i},f_{j}]=n_{j}^{T}G^{T}Qg_{i}-n_{i}^{T}G^{T}Qn_{j}\quad i,j=1,\dots,m,$ (3.4)
@@
The result is a direct consequence of Theorem A.1 in the Appendix A. ∎
An optimality condition for linear pH systems has been studied in [39]. In order to compare our results with those in [39], we see that the optimality condition (3.5) is automatically satisfied in their framework, since there $N=I_{m}$ is a standing hypothesis. Additionally, observe that we do not need the normality hypothesis used in [39], since our problem (3.2) does not involve terminal constraints, while in [39] the terminal point is fixed.
@@
$[f_{0},f_{i}]=\begin{bmatrix}-Ag_{i}\\ n_{i}^{T}G^{T}QAx-x^{T}QG(M+M^{T})G^{T}Qg_{i}\end{bmatrix},$
@@
$d_{i}$ $=$ $\begin{bmatrix}p\\ 1\end{bmatrix}^{T}\left[f_{0},[f_{0},f_{i}]\right]$
$=$ $p^{T}A^{2}g_{i}+n_{i}^{T}G^{T}QA^{2}x-2g_{i}^{T}QGMG^{T}QAx+2x^{T}QGMG^{T}QAg_{i}.$
@@
$\begin{bmatrix}p\\ 1\end{bmatrix}^{T}\left[f_{j},[f_{0},f_{i}]\right]=n_{i}^{T}G^{T}QAG_{j}-2g_{i}^{T}QGMG^{T}Qg_{j}+n_{j}^{T}G^{T}QAg_{i}.$
@@
which is positive semidefinite in view of the standing hypotheses for $Q$ and $R$. Moreover, whenever $R$ is definite, so will be $W,$ leading then to a feedback expression for the singular control from (2.7). In the latter case, along such trajectories, the solution is asymptotically stable since the dissipation term acts in this direction. Additionally, under the hypotheses used in [39], namely, (3.7) together with
@@
the expression of $W$ also coincides with (3.8), which is negative definite in view of assumption (3.9). Here we see directly that under the conditions (3.7) and (3.9) (as in [39]), a feedback solution exists. On the other hand our analysis is more general, since we get analogous results under noticeably weaker assumptions.
@@
In Theorems 3.1 and 3.2, we have derived the Goh and generalized Legendre-Clebsch conditions for linear port-Hamiltonian (pH) systems, which correspond to second-order optimality conditions. Higher-order conditions (see, e.g., [16, 18]) involve increasingly complex higher-order Lie brackets. Even for linear pH systems, these expressions become overly intricate, limiting their practical utility and insight.
@@
To illustrate the elegance of the optimality conditions, consider the following example of a second order linear control problem that is discussed in detail in [10] and arises e.g. in the control of high rise buildings [44]:
@@
with real symmetric positive definite matrices $M,K$ and $B$ of full column rank. In contrast to [10] we assume that the damping does not work on the whole position vector, i.e., that $D$ is only positive semidefinite. Setting $x=\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix}=\begin{bmatrix}M\dot{q}\\ q\end{bmatrix}$, $R=\begin{bmatrix}D&0\\ 0&0\end{bmatrix}$, $J=\begin{bmatrix}0&-I\\ I&0\end{bmatrix}$, $Q=\begin{bmatrix}M^{-1}&0\\ 0&K\end{bmatrix}$, $G=\begin{bmatrix}B\\ 0\end{bmatrix}$, and introducing a collocated output $y=G^{T}Qx$, the system has the form (3.1). Unfortunately the condition ${\rm Im}(G)\cap{\rm Ker}(RQ)=\{0\}$ that is used in [39] does not hold, since $DM^{-1}$ has a kernel and $W=-2B^{T}M^{-1}DM^{-1}B$ is only semidefinite. Hence the techniques in  [39] cannot be applied. Also the linear system with $W$ for the feedback is, in general, not (uniquely) solvable. However, still a (non-unique) feedback solution may exist if the right hand side $d$ is in the image of $W$, see [10, 27]. In case $M$ is only semidefinite then one can rewrite the system in descriptor form as in (1.5) and the presented procedures apply to this case. We then obtain the optimality conditions as in Section 2.3.
In this subsection we have seen that by using the general theory we can generalize the conditions for optimality presented in [39]. In the next subsection we consider the case of control bounds.
@@ 3.2 Optimality conditions for linear pH problems subject to control bounds
$s_{i}\coloneqq\mathcal{H}_{u_{i}}=\begin{bmatrix}p\\ 1\end{bmatrix}^{T}f_{i}=p^{T}g_{i}+x^{T}QGn_{i}+\ell_{i}.$ (3.15)
@@
In view of the maximum condition of Pontryagin’s Maximum Principle [33], one then has that $W_{i}$ is contained, up to a set of measure zero, in $Z_{i}$.
Now, following the notation in [39, Theorem 8], for any subset $\mathcal{I}\subseteq\{1,\dots,m\}$, we set
@@
$\left\{\begin{array}[]{rl}\displaystyle\frac{d}{dt}p^{T}&=-p^{T}A-p_{n+1}\left(x^{T}QG(Y+Y^{T})G^{T}Q+u^{T}N^{T}G^{T}Q\right),\\[4.30554pt] \dot{p}_{n+1}&=0,\end{array}\right.$
recalling that $p_{n+1}\equiv 1,$ since the involved terminal cost is $x_{n+1}(T).$
@@
$\begin{bmatrix}p\\ 1\end{bmatrix}^{T}[f_{i},f_{j}]=n_{j}^{T}G^{T}Qg_{i}-n_{i}^{T}G^{T}Qn_{j}=0,$ (3.16)
on any open interval $(a,b)$ on which
@@
The proof follows from Theorems A.3 and A.4 in the Appendix. ∎
@@
When $N=0$ or $N=I_{m}$, using the notation introduced in (2.3) for the vector fields $f_{i},$ one has that $[f_{i},f_{j}]=0$ in the whole space $\mathbb{R}^{n},$ independently of the trajectories. Therefore, the Goh conditions hold trivially. Additionally, for $N=0,$ the matrix in the generalized Legendre-Clebsch condition of Theorem 3.4 reads
$-2G_{\mathcal{I}}^{T}QGYG^{T}QG_{\mathcal{I}},$ (3.18)
@@
$0=p^{T}A^{2}G_{\mathcal{I}}$
$+x^{T}\left[2QGYG^{T}QAG_{\mathcal{I}}-2A^{T}QGYG^{T}QG_{\mathcal{I}}+A^{T}A^{T}QGN_{\mathcal{I}}\right]$
@@
00 0 $=$ $p^{T}A^{2}G_{\mathcal{I}}+x^{T}\left[2QGYG^{T}QAG_{\mathcal{I}}-2A^{T}QGYG^{T}QG_{\mathcal{I}}\right]$
@@
00 0 $=$ $p^{T}A^{2}G_{\mathcal{I}}+x^{T}\left[2QGYG^{T}Q(J-R)-2A^{T}QGYG^{T}+A^{T}A^{T}\right]QG_{\mathcal{I}}$
@@
$s_{\mathcal{I}}=p^{T}G_{\mathcal{I}}+p_{n+1}\left(x^{T}QGN_{\mathcal{I}}+\ell_{\mathcal{I}}\right).$
@@
$\dot{s}_{\mathcal{I}}$ $=$ $\frac{d}{dt}p^{T}G_{\mathcal{I}}+p_{n+1}\dot{x}^{T}QGN_{\mathcal{I}}$
$=$ $-p^{T}AG_{\mathcal{I}}-p_{n+1}x^{T}QG(Y+Y^{T})G^{T}QG_{\mathcal{I}}$
$-u^{T}\overbrace{N^{T}G^{T}QG_{\mathcal{I}}}^{(N^{T}G^{T}QG)_{\mathcal{I}}}+x^{T}A^{T}QGN_{\mathcal{I}}+p_{n+1}u^{T}\underbrace{G^{T}QGN_{\mathcal{I}}}_{(G^{T}QGN_{\mathcal{I})}}.$
@@
$\dot{s}_{\mathcal{I}}=-p^{T}AG_{\mathcal{I}}-x^{T}QG(Y+Y^{T})G^{T}QG_{\mathcal{I}}+x^{T}A^{T}QGN_{\mathcal{I}}.$ (3.21)
@@
$\ddot{s}_{\mathcal{I}}=$ $p^{T}A^{2}G_{\mathcal{I}}x^{T}\left[2QGYG^{T}QAG_{\mathcal{I}}-2A^{T}QGYG^{T}QG_{\mathcal{I}}+A^{T}A^{T}QGN_{\mathcal{I}}\right]$ (3.22)
@@
$\ddot{s}_{\mathcal{I}}=$ $p^{T}A^{2}G_{\mathcal{I}}+x^{T}\left[2QGYG^{T}QAG_{\mathcal{I}}-2A^{T}QGYG^{T}QG_{\mathcal{I}}\right]$
@@
$\ddot{s}_{\mathcal{I}}=$ $p^{T}A^{2}G_{\mathcal{I}}+x^{T}\left[2QGYG^{T}Q(J-R)-2A^{T}QGYG^{T}+A^{T}A^{T}\right]QG_{\mathcal{I}}$
@@
where we replaced the matrix $A$ by its expression in (3.3).

### v3

@@ 3.2 Optimality conditions for nonlinear ordinary pH systems with control bounds
with real numbers $\underline{u}_{i}\leq\overline{u}_{i},$ for each $i=1,\dots,m$. In this case, the Goh optimality condition is restricted to the components that lie in the interior of the control set and one has the following optimality condition.
@@
for some pair $i,j=1,\ldots,m,$ the Goh conditions (3.4) must necessarily hold.
The proof of Theorem 3.4 is not a straightforward consequence of the unconstrained version of Theorem 2.1. For more details, the reader is referred to [ABDL12, FraTon13].
@@
Under the hypotheses of Theorem 3.4, let $(a,b)$ be an open interval on which
$\left\{\begin{array}[]{ll}\underline{u}_{i}<u^{*}_{i}(t)<\overline{u}_{i},\quad&\text{for }i\in S,\\ u^{*}_{i}(t)=\underline{u}_{i},\quad&\text{for }i\in\underline{B},\\ u^{*}_{i}(t)=\overline{u}_{i},\quad&\text{for }i\in\overline{B}.\end{array}\right$.
Thus, $\{S,\underline{B},\overline{B}\}$ is a partition of $\{1,\dots,m\}$ into singular, lower-bang and upper-bang components of the control $u^{*}.$ One gets, for $i\in S$, that
$\lambda^{T}[f_{0},[f_{0},f_{i}]]+\sum_{j\in S}u_{j}^{*}\lambda^{T}[f_{j},[f_{0},f_{i}]]$ (3.9)
$+\sum_{j\in\underline{B}}\underline{u}_{j}\lambda^{T}[f_{j},[f_{0},f_{i}]]+\sum_{j\in\overline{B}}\overline{u}_{j}\lambda^{T}[f_{j},[f_{0},f_{i}]]=0.$
holds on $(a,b)$. Additionally, the $|S|\times|S|$-matrix $W_{S}(t)$ with entries $w_{ij}=\lambda^{T}[f_{j},[f_{0},f_{i}]],$ for $i,j\in S$, is negative semidefinite.
@@
The expression (3.9) follows from calculations analogous to (2.6)–(2.7), (2.9) and Goh condition of Theorem 3.4. The negative semidefiniteness of $W_{S}$ holds in view of [FraTon13, Theorem 4.1]. ∎
Remark 3.3 (A feedback formula for the singular controls).
The matrix $W_{S}$ in Theorem 3.5 is the submatrix of $W$ in (2.11) corresponding to the singular components of the control. Whenever $W_{S}$ is negative definite, by using expression (3.9), one can write the singular control in terms of the state variable $x$, the adjoint variable $\lambda$ and the bang controls.
@@ 3.3 Optimality conditions for nonlinear pH descriptor systems
In this section we extend the results from the previous two subsections to a special class of pH descriptor systems. For general descriptor systems, optimality conditions have been derived in [KunMeh08], see [KunM24] for a detailed exposition. But again these do not exploit the particular structure.
It has been shown in [CamKM12] that by an appropriate regularization procedure, using transformations, derivatives, and feedback designs, a general descriptor system can always be transformed to a so-called index-one or strangeness-free system, see [KunM24]. Then, for a system of the form (1.1)–(1.3), after appropriate coordinate changes, we may assume that the system has the structure
@@
$\left(\begin{bmatrix}J_{11}&J_{12}\\ -J_{12}^{T}&J_{22}\end{bmatrix}(x)-\begin{bmatrix}R_{11}&R_{12}\\ R_{12}^{T}&R_{22}\end{bmatrix}(x)\right)\begin{bmatrix}Q_{11}&Q_{12}\\ Q_{21}&Q_{22}\end{bmatrix}\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix}$
@@
Observe that the condition $Q^{T}E=E^{T}Q\geq 0$ implies that $Q_{12}=Q_{21}^{T}=0$ and $Q_{11}\geq 0$. In the following, to keep the presentation simple, we also assume that $G_{2}(x)=0$. If this is not the case, then the following reduction procedure will lead to a system with a feed-through term for which an extra transformation is necessary, see [MehU23]. The assumption that the system is of index at most one means that from the algebraic equation
$0=$ $[(J_{22}(x_{1},x_{2})-R_{22}(x_{1},x_{2}))Q_{21}+(-J_{12}^{T}(x_{1},x_{2})-R_{12}^{T}(x_{1},x_{2}))Q_{11}x_{1}$
$+[(J_{22}(x_{1},x_{2})-R_{22}(x_{1},x_{2}))Q_{22}x_{2},$
the variable $x_{2}$ can be expressed via the Implicit Function Theorem as a function of $x_{1}$ and inserted in the coefficients. Due to the structure of $G(x)$, $x_{2}$ contributes to the cost functional only indirectly as argument of $G$. Thus, we can first solve the optimal control problem for $x_{1}$ and obtain $x_{2}$ in a post-processing step. This means that the dynamics is described by the ordinary pH system
@@
4 Optimality conditions for linear pH systems
@@ 4.1 Optimality conditions for linear pH systems without control constraints
  • $J\in\mathbb{R}^{n,n}$ is skew-symmetric;
  • $R\in\mathbb{R}^{n,n}$ is symmetric positive semidefinite;
  • $Q\in\mathbb{R}^{n,n}$ is symmetric positive definite;
  • $G=\begin{bmatrix}g_{1}&\ldots&g_{m}\end{bmatrix}\in\mathbb{R}^{n,m}$ has full rank $m\leq n;$
@@
Note that more general linear port-Hamiltonian formulations are possible, we follow here the formulation considered in [FauKMPSW25, FauMPSW22, SchPFWM23].
@@
  • $Y\in\mathbb{R}^{n,n}$ is symmetric and positive semidefinite;
  • $N=\begin{bmatrix}n_{1}&\ldots&n_{m}\end{bmatrix}\in\mathbb{R}^{m,m}$;
@@
$f_{i},f_{j}]=\begin{bmatrix}0\\ \vdots\\ 0\\ n_{j}^{T}G^{T}Qg_{i}-n_{i}^{T}G^{T}Qg_{j}\end{bmatrix},\quad i,j=1,\dots,m,$
@@
$\begin{bmatrix}\lambda\\ 1\end{bmatrix}^{T}[f_{i},f_{j}]=n_{j}^{T}G^{T}Qg_{i}-n_{i}^{T}G^{T}Qn_{j},\quad i,j=1,\dots,m,$ (4.4)
@@
The result follows from Theorem 2.1. ∎
An optimality condition for linear pH systems has been studied in [SchPFWM23]. In order to compare our results with those in [SchPFWM23], we see that the optimality condition (4.5) is automatically satisfied in their framework, since there $N=I_{m}$ is a standing hypothesis. Observe that, while in [SchPFWM23] the terminal point is fixed—requiring the imposition of a normality condition—our problem (4.2) does not include terminal constraints. Although this may be seen as a limitation in certain contexts, it has the advantage of ensuring normality without additional assumptions. The result also generalized the tracking control result discussed in [SchZBSTW24].
@@
$f_{0},f_{i}]=\begin{bmatrix}-Ag_{i}\\ n_{i}^{T}G^{T}QAx-x^{T}QG(M+M^{T})G^{T}Qg_{i}\end{bmatrix}$,
@@
$d_{i}$ $=$ $\begin{bmatrix}\lambda\\ 1\end{bmatrix}^{T}\left[f_{0},[f_{0},f_{i}]\right]$
$=$ $\lambda^{T}A^{2}g_{i}+n_{i}^{T}G^{T}QA^{2}x-2g_{i}^{T}QGMG^{T}QAx+2x^{T}QGMG^{T}QAg_{i}.$
@@
$\begin{bmatrix}\lambda\\ 1\end{bmatrix}^{T}\left[f_{j},[f_{0},f_{i}]\right]=n_{i}^{T}G^{T}QAG_{j}-2g_{i}^{T}QGMG^{T}Qg_{j}+n_{j}^{T}G^{T}QAg_{i}.$
@@
which is positive semidefinite, in view of the standing hypotheses for $Q$ and $R$. Moreover, whenever $R$ is definite, so will be $W,$ leading then to a feedback expression for the singular control from (2.11). In the latter case, along such trajectories, the solution is asymptotically stable since the dissipation term acts in this direction. Additionally, under the hypotheses used in [SchPFWM23], namely, (4.7) together with
@@
the expression of $W$ also coincides with (4.8), which is negative definite in view of assumption (4.9). Here we see directly that under the conditions (4.7) and (4.9) (as in [SchPFWM23]), a feedback solution exists. On the other hand our analysis is more general, since we get analogous results under noticeably weaker assumptions.
@@
In Theorems 4.1 and 4.2, we have derived the Goh and generalized Legendre-Clebsch conditions for linear port-Hamiltonian (pH) systems, which correspond to second-order optimality conditions. The singular control is characterized when the singularity is of order 1 (i.e., the control appears explicitly in the derivative $\frac{d^{2}}{dt^{2}}\mathcal{H}_{u}$). Higher-order conditions (see, e.g., [Knobloch80, Krener1977]) involve increasingly complex higher-order Lie brackets. Even for linear pH systems, these expressions become overly intricate, limiting their practical utility and insight.
@@
To illustrate the elegance of the optimality conditions, consider the following example of a second order linear control problem that is discussed in detail in [FauKMPSW25] and arises e.g. in the control of high rise buildings [WarsBohmSawoTari21]:
@@
with real symmetric positive definite matrices $M,K$ and $B$ of full column rank. In contrast to [FauKMPSW25] we assume that the damping does not work on the whole position vector, i.e., that $D$ is only positive semidefinite. Setting $x=\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix}=\begin{bmatrix}M\dot{q}\\ q\end{bmatrix}$, $R=\begin{bmatrix}D&0\\ 0&0\end{bmatrix}$, $J=\begin{bmatrix}0&-I\\ I&0\end{bmatrix}$, $Q=\begin{bmatrix}M^{-1}&0\\ 0&K\end{bmatrix}$, $G=\begin{bmatrix}B\\ 0\end{bmatrix}$, and introducing a collocated output $y=G^{T}Qx$, the system has the form (4.1). Unfortunately, the condition ${\rm Im}(G)\cap{\rm Ker}(RQ)=\{0\}$ that is used in [SchPFWM23] does not hold, since $DM^{-1}$ has a kernel and $W=-2B^{T}M^{-1}DM^{-1}B$ is only semidefinite. Hence the techniques in  [SchPFWM23] cannot be applied. Also the linear system with $W$ for the feedback is, in general, not (uniquely) solvable. However, still a (non-unique) feedback solution may exist if the right hand side $d$ is in the image of $W$, see [FauKMPSW25, Meh91]. In case $M$ is only semidefinite then one can rewrite the system in descriptor form as in (1.4) and the presented procedures apply to this case. We then obtain the optimality conditions as in Section 3.3.
In this subsection we have seen that, by using the general theory, we can generalize the conditions for optimality presented in [SchPFWM23]. In the next subsection we consider the case of control bounds.
@@ 4.2 Optimality conditions for linear pH problems subject to control bounds
$s_{i}\coloneqq\mathcal{H}_{u_{i}}=\begin{bmatrix}\lambda\\ 1\end{bmatrix}^{T}f_{i}=\lambda^{T}g_{i}+x^{T}QGn_{i}+\ell_{i}.$ (4.15)
@@
In view of the maximum condition of Pontryagin’s Maximum Principle [Pontryagin], one then has that $W_{i}$ is contained, up to a set of measure zero, in $Z_{i}$.
Now, following the notation in [SchPFWM23, Theorem 8], for any subset $\mathcal{I}\subseteq\{1,\dots,m\}$, we set
@@
$\left\{\begin{array}[]{rl}\displaystyle\frac{d}{dt}\lambda^{T}&=-\lambda^{T}A=-\lambda_{n+1}\left(x^{T}QG(Y+Y^{T})G^{T}Q+u^{T}N^{T}G^{T}Q\right),\\[4.30554pt] \dot{\lambda}_{n+1}&=0,\end{array}\right$.
recalling that $\lambda_{n+1}\equiv 1,$ since the involved terminal cost is $x_{n+1}(T).$
@@
$\begin{bmatrix}\lambda\\ 1\end{bmatrix}^{T}[f_{i},f_{j}]=n_{j}^{T}G^{T}Qg_{i}-n_{i}^{T}G^{T}Qn_{j}=0,$ (4.16)
on any open subinterval $(a,b)$ of $[0,T]$ on which
@@
The proof follows from Theorems 3.4 and 3.5. ∎
@@
When $N=0$ or $N=I_{m}$, using the notation introduced in (3.3) for the vector fields $f_{i},$ one has that $f_{i},f_{j}]=0$ in the whole space $\mathbb{R}^{n},$ independently of the trajectories. Therefore, the Goh conditions hold trivially. Additionally, for $N=0,$ the matrix in the generalized Legendre-Clebsch condition of Theorem 4.4 reads
$2G_{\mathcal{I}}^{T}QGYG^{T}QG_{\mathcal{I}},$ (4.18)
@@
$0=\lambda^{T}A^{2}G_{\mathcal{I}}+x^{T}\left[2QGYG^{T}QAG_{\mathcal{I}}-2A^{T}QGYG^{T}QG_{\mathcal{I}}+A^{T}A^{T}QGN_{\mathcal{I}}\right]$
@@
$0$ $=$ $\lambda^{T}A^{2}G_{\mathcal{I}}+x^{T}\left[2QGYG^{T}QAG_{\mathcal{I}}-2A^{T}QGYG^{T}QG_{\mathcal{I}}\right]$
@@
$0$ $=$ $\lambda^{T}A^{2}G_{\mathcal{I}}$
$+$ $x^{T}\left[2QGYG^{T}Q(J-R)-2A^{T}QGYG^{T}+A^{T}A^{T}\right]QG_{\mathcal{I}}$
@@
$s_{\mathcal{I}}=\lambda^{T}G_{\mathcal{I}}+\lambda_{n+1}\left(x^{T}QGN_{\mathcal{I}}+\ell_{\mathcal{I}}\right).$
@@
$\dot{s}_{\mathcal{I}}$ $=$ $\frac{d}{dt}\lambda^{T}G_{\mathcal{I}}+\lambda_{n+1}\dot{x}^{T}QGN_{\mathcal{I}}$
$=$ $-\lambda^{T}AG_{\mathcal{I}}-\lambda_{n+1}x^{T}QG(Y+Y^{T})G^{T}QG_{\mathcal{I}}$
$-u^{T}\underbrace{N^{T}G^{T}QG_{\mathcal{I}}}_{(N^{T}G^{T}QG)_{\mathcal{I}}}+x^{T}A^{T}QGN_{\mathcal{I}}+\lambda_{n+1}u^{T}\underbrace{G^{T}QGN_{\mathcal{I}}}_{(G^{T}QGN)_{\mathcal{I}}}.$
@@
$\dot{s}_{\mathcal{I}}=-\lambda^{T}AG_{\mathcal{I}}-x^{T}QG(Y+Y^{T})G^{T}QG_{\mathcal{I}}+x^{T}A^{T}QGN_{\mathcal{I}}.$ (4.21)
@@
$\ddot{s}_{\mathcal{I}}$ $=\lambda^{T}A^{2}G_{\mathcal{I}}+x^{T}\left[2QGYG^{T}QAG_{\mathcal{I}}-2A^{T}QGYG^{T}QG_{\mathcal{I}}+A^{T}A^{T}QGN_{\mathcal{I}}\right]$ (4.22)
@@
$\ddot{s}_{\mathcal{I}}=$ $\lambda^{T}A^{2}G_{\mathcal{I}}+x^{T}\left[2QGYG^{T}QAG_{\mathcal{I}}-2A^{T}QGYG^{T}QG_{\mathcal{I}}\right]$
@@
$\ddot{s}_{\mathcal{I}}=$ $\lambda^{T}A^{2}G_{\mathcal{I}}+x^{T}\left[2QGYG^{T}Q(J-R)-2A^{T}QGYG^{T}+A^{T}A^{T}\right]QG_{\mathcal{I}}$
@@
where we replaced the matrix AA by its expression in (4.3).